length = 6.5 m
width = ?
heigth = 2.5 m
volume = 65 yd to the third power
what is the width?

Answer by checkley75(3666)

L*H*W=VOLUME
6.5*2.5*X=65
16.25X=65
X=65/16.25
X=4 FOR THE WIDTH.